Apply Gauss’s Divergence theorem to evaluate taken over the

sphere being the direction cosines of the external normal to the

sphere.


1
Expert's answer
2021-10-27T16:25:40-0400

Here the given question is not complete and it is not from the related subject. So, I am answering the approach of the above question.

Gauss Divergence Theorem:

The surface integral of a vector A over the closed surface = Volume integral of the divergence of a vector field A over the volume enclosed by closed surface.


"\Rightarrow \oiint_s \overrightarrow{A}.\overrightarrow{dS}=\iiint_v(\overrightarrow{{\nabla}}.\overrightarrow{A})dV"


Let there is a surface S, which encloses a surface A. Let A be the vector field in the given region.

Let the volume of the small part of the sphere "\Delta V_j" which is bounded by a surface "S_j"


"\Rightarrow \oiint_s \overrightarrow{A}.\overrightarrow{dS}"


Now, we will integrate the volume,


"\Rightarrow \Sigma \oiint_{sj} \overrightarrow{A}.\overrightarrow{dS_j}=\oiint_s \overrightarrow{A}.\overrightarrow{dS}...(i)"


Now, in the above equation, multiply and divide by the "\Delta V_i"


"\Rightarrow \oiint_s \overrightarrow{A}.\overrightarrow{dS}=\Sigma{\frac{1}{\Delta{V_i}}}(\oiint\overrightarrow{A}\overrightarrow{dS_j})\Delta{V_i}"


Let volume is divide into the infinite elementary volume,


"\Rightarrow \oiint_s \overrightarrow{A}.\overrightarrow{dS}=\lim_{\Delta v_i\rightarrow 0}\Sigma{\frac{1}{\Delta{V_i}}}(\iint\overrightarrow{A}\overrightarrow{dS_j})\Delta{V_i}...(ii)"


"\lim_{\Delta v_i\rightarrow 0}[\Sigma{\frac{1}{\Delta{V_i}}}(\oiint\overrightarrow{A}\overrightarrow{dS_i})]=(\overrightarrow{V}.\overrightarrow{A})"


"\oiint{(\overrightarrow{A}.\overrightarrow{dS})}=\Sigma{(\overrightarrow{\nabla}.\overrightarrow{A})}\Delta{V}_i"

Hence,

"\oiint{(\overrightarrow{A}.\overrightarrow{dS})}=\iiint_v(\overrightarrow{\nabla}.\overrightarrow{A})dV"